# Integrating Random Regret Minimization-Based Discrete Choice Models with Mixed Integer Linear Programming for Revenue Optimization

### v2

@@ Abstract
This paper explores the critical domain of Revenue Management (RM) within Operations Research (OR), focusing on intricate pricing dynamics. Utilizing Mixed Integer Linear Programming (MILP) models, the study enhances revenue optimization by considering product prices as decision variables and emphasizing the interplay between demand and supply. Recent advancements in Discrete Choice Models (DCMs), particularly those rooted in Random Regret Minimization (RRM) theory, are investigated as potent alternatives to established Random Utility Maximization (RUM) based DCMs. Despite the widespread use of DCMs in RM, a significant gap exists between cutting-edge RRM-based models and their practical integration into RM strategies. The study addresses this gap by incorporating an advanced RRM-based DCM into MILP models, addressing pricing challenges in both capacitated and uncapacitated supply scenarios. The developed models demonstrate feasibility and offer diverse interpretations of consumer choice behavior, drawing inspiration from established RUM-based frameworks. This research contributes to bridging the existing gap in the application of advanced DCMs within practical RM implementations.
@@ 1 Introduction
Advancements in Operations Research (OR) Modeling, coupled with innovative solving methods, have efficiently addressed large-scale challenges across economics, transportation, supply chain, revenue management, etc. The primary goal of these models is to optimize revenue, minimize costs, or mitigate pollution. Consequently, sophisticated demand models are under development to precisely quantify these factors. Recognizing the intricate interplay between demand and supply, our approach integrates both aspects, emphasizing simultaneous consideration of the supply side during supply chain planning. Notably, the pricing decision involves suppliers setting prices while buyers determine their willingness to pay, exemplified by the varying affordability of a \$10 rose in different neighborhoods. This underscores the crucial need for sellers to comprehend and anticipate customer purchasing behavior. Concurrently, Revenue Management (RM), within the realm of OR, concentrates on maximizing revenue through strategic pricing and capacity allocation. Our work uniquely combines economic and OR perspectives, incorporating customer behavior modeling for optimized supply systems.
Historically, demand and revenue-maximizing prices were determined using simple linear demand functions. However, this approach lacks accuracy due to individual variations in purchase preferences and willingness to pay. RM addresses these challenges by predicting customer behavior to enhance revenue talluri2006theory .
@@
An essential assumption in RM is the variability of product prices for each customer. Precise understanding of customer behavior is necessary to optimize revenue. Discrete Choice Models (DCMs), which elucidate and predict customer choices, significantly contribute to customer behavior modeling. These models consider customer preferences, leading to efficient resource utilization and substantial revenue improvements. Numerous empirical studies underscore the impact of discrete choice modeling in diverse sectors such as RM talluri2004revenue , labor market kornstad2007discrete , food industry gracia2008demand , and airline industry vulcano2010om .
@@
Random Utility Maximization Theory (RUM), established by mcfadden1978modelling ; mcfadden2000mixed , forms the foundation for discrete choice modeling. The core assumption of RUM is that decision-makers are perfectly rational and aim to maximize their utilities. Recent studies, such as those by chorus2008random ; chorus2010new ; chorus2012rrm propose an alternative approach by investigating DCMs derived from Random Regret Minimization (RRM) where the customer tries to minimize the regret caused by not choosing an alternative. Conditions exist where DCMs derived from either theory outperform the other, revealing diverse customer preferences hensher2013random ; greene2011ru ; li2018hybrid .
@@
While existing literature often examines demand and supply separately or investigates supply based on a simplistic demand formulation, a comprehensive approach is needed for effective revenue generation pacheco2016new . Neglecting the interplay between supply and demand, or focusing solely on one side, can lead to suboptimal pricing strategies, resulting in reduced gains, the bullwhip effect, and discordance in supply chains. Recognizing that customer decisions drive demand, it underscores the vital importance of optimizing and coordinating both supply and demand concurrently. In this regard, pacheco2017integrating innovatively crafted a framework that characterizes how demand responds to suppliers’ decisions, linking supply and demand through price as a crucial variable. They linearized DCMs into predictors of demand using a Mixed Integer Linear Programming (MILP) model, emphasizing price variables in the objective function.
@@
In related literature, other works have incorporated DCMs in OR to forecast disaggregated demand and maximize revenue. For example, pacheco2016new ; paneque2017integrating addressed challenges arising from the non-linearity of DCM models in optimization. To address this, utility functions (RUM-based) were integrated into MILP constraints, introducing individual utility functions as scores to be maximized in discrete optimization, crucial for characterizing the demand side of the OR problem.
Existing literature in RM predominantly estimates demand based on RUM theory. However, RRM theory, a recent challenger to RUM, has garnered significant attention among researchers. To our knowledge, there is currently no choice-based optimization method grounded in RRM. Creating a framework akin to pacheco2017integrating but rooted in RRM would empower researchers to evaluate both demand and supply concurrently, offering a distinct interpretation of customer behavior.
In this paper, a novel MILP model based on RRM and the work of pacheco2016new ; pacheco2017integrating has been introduced in the realm of choice-based optimization. This framework serves to connect economic theories with the optimization of supply chains.
@@ 2 Related Work
Within RM, pricing decisions hold prominent importance among various aspects to increase the revenue. Several RM techniques aim to leverage customer heterogeneity for a product set at a specific time. They often focus on fixing alternatives or products and time dimensions while exploring customer differences. In our context, we emphasize the significance of customer heterogeneity in purchasing behavior. If all customers perceive a product uniformly with identical purchasing behavior, the customer dimension diminishes, leading to unexploited variations in willingness to pay. These variations, influenced by socio-economic factors like marital status, salary, age, gender, etc., play a critical role in maximizing revenue. On the other hand, optimizing revenue hinges on accurate demand prediction, crucial for effective supply planning and pricing strategies. The demand function is influenced by various variables, encompassing customer heterogeneity, prices, etc talluri2006theory ; SADRALASHRAFI201882 . Specifically, researchers have extensively focused on DCMs, particularly due to their capacity to predict demand by accounting for consumer heterogeneity in tastes and attributes talluri2006theory ; mcfadden2000mixed ; chorus2008random . RUM-based DCMs assume decision-makers are rational utility maximizers, and models derived from these theories aim to reveal influential factors shaping customer choice preferences (e.g., see kemperman2021review for a recent review). However, in some cases, customers prefer regret minimization over utility maximization when making choices. Therefore, deriving DCMs from suitable behavioral theories is crucial hensher2013random . In this regard, chorus2014random introduced a regret-based discrete choice model for consumer preference analysis with multinomial choice sets. Their empirical assessment, based on 43 comparisons between RUM and RRM models in the literature, revealed the superiority of RRM-based DCMs over RUM-based ones. The study demonstrated that both RRM-based and hybrid RRM-RUM-based DCMs consistently outperformed their RUM-based counterparts in terms of prediction accuracy and overall model fit. The authors emphasized that although performance differences were minor, the models led to significant managerial implications and diverse interpretations of behavior.
@@
The reference masiero2019understanding conducted a choice experiment to investigate tourists’ hotel preferences. They created hypothetical hotel alternatives and randomized factors, including price, hotel rating, and walking time to key attractions, to form choice sets. RRM and RUM-based DCMs were utilized to estimate the coefficients of these factors where they figured out the RRM-based DCMs performed significantly better than their RUM-based counterparts. Additionally, they showed that quantitative simulations can be performed to study the demand and pricing issues while incorporating customers’ heterogeneity into the model.
The reference andersson1998passenger made a pioneering contribution to Revenue Management by focusing on modeling passenger preferences using DCMs. Following andersson1998passenger , gallego2004managing addressed a general question in RM related to consumer choice behavior (RUM). They explored the network revenue management problem with flexible products, considering scenarios where demand was either exogenously given or generated by a consumer choice model based on RUM-based DCM. Their findings indicated that the optimal value of the stochastic optimization problem closely approximated that of a deterministic linear program in both settings.
The reference liu2008choice delved into the implications of the concept introduced by gallego2004managing in RM. They developed a Choice-based Deterministic Linear Programming (CDLP) model under consumer choice behavior, focusing on providing suitable offer sets, seats, and flights in the airline industry while adhering to capacity constraints over time. Their study demonstrated that the CDLP, inspired by gallego2004managing , serves as a deterministic approximation of the stochastic formulation, with its solution being an asymptotic optimal solution for the original stochastic problem influenced by the stochasticity of customer choice probability.
The reference bront2009column expanded on the CDLP introduced by liu2008choice . They addressed the issue of overlapping offer sets in the airline industry, where airlines typically present affordable morning tickets in one set and pricier afternoon tickets in another set. This extension considered combinations such as expensive and morning tickets, providing insights into more realistic scenarios. They applied column generation method to tackle the dimensionaliy issue of the problem. Particularly, meta-heauristc methods seem to be a promising approach in approximating high quality solutions for large-sized problems (e.g., see Haeri_Boroujeni_2023 ; mehrabi2023efficient ; boroujeni2021data ; 9721496 ; https://doi.org/10.13140/rg.2.2.29416.03849 ).
The reference azadeh2015impact introduced a modified CDLP model, solving it on a synthetic dataset to obtain optimal offer sets and an upper bound on revenue. Additionally, they suggested a non-parametric model for estimating choice probabilities, utilizing the synthetic data employed in the CDLP. References bierlaire2016demand ; PACHECOPANEQUE202126 ; PACHECOPANEQUE2022105985 introduced the linearized formulation of demand based on customer choice behavior, employing a multinomial logit model, a type of DCM which assumes the unobserved utility of the customer is independent of the attributes of the customer, and acknowledging the non-linearity and non-convexity of the resulting model. They integrated it into a MILP model, marking the emergence of RUM-based MILP models in RM. Subsequently, pacheco2017integrating linearized a mixed multinomial logit DCM model, a model where the unobserved utility of the customer is deemed to be correlated with the attributes of the customer, through MILP, aiming to maximize revenue by simultaneously exploring supply (price and capacity) and demand (price, utility levels, customer satisfaction) factors.
The reference paneque2019passenger delved into an economics-oriented approach, emphasizing passenger satisfaction maximization by integrating DCMs into MILP models. The objective was to enhance the public transportation network by utilizing toll revenue from highways, ensuring a balance between economic considerations and passenger contentment.
@@
Generally, machine algorithms can be broadly categorized into one of the following three types: supervised, unsupervised, and agent-based learning. Paper boroujeni2024comprehensive provides a comprehensive overview of the existing ML techniques along with their types, characteristics and potential applications. Indeed, machine learning (ML) tools such as MNL and MMNL are used to predict customers’ choices based on RUM theory. For a recent literature on this regard, see hillel2021systematic ; soleymani_forecasting_2024 ; boroujeni2024ic ; 9814765 ; yousefpour2023unsupervised . Reinforcement learning as an online learning method can be also leveraged to help RM. For a recent literature on this regard, see jebellat2023reinforcement ; ran2024development ; amiri2023rank ; amirimargavi2023lowrank .
@@
In conclusion, DCMs are state-of-the-art for estimating dis-aggregate demand, showing promise in RM. The literature demonstrates integrating DCMs into optimization models for better supply-demand interactions. However, a gap exists between DCMs derived from theories like RRM and optimization formulations. Finally, pacheco2016new ; pacheco2017integrating ; PACHECOPANEQUE202126 ; PACHECOPANEQUE2022105985 developed a RUM-based DCM framework where customers are utility maximizers. We extended the framework for RRM-based DCM introduced by chorus2010new , where the customers are regret minimizers.

### v3

@@ Abstract
This paper explores the critical domain of Revenue Management (RM) within Operations Research (OR), focusing on intricate pricing dynamics. Utilizing Mixed Integer Linear Programming (MILP) models, the study enhances revenue optimization by considering product prices as decision variables and emphasizing the interplay between demand and supply. Recent advancements in Discrete Choice Models (DCMs), particularly those rooted in Random Regret Minimization (RRM) theory, are investigated as potent alternatives to established Random Utility Maximization (RUM) based DCMs. Despite the widespread use of DCMs in RM, a significant gap exists between cutting-edge RRM-based models and their practical integration into RM strategies. The study addresses this gap by incorporating an advanced RRM-based DCM into MILP models, addressing pricing challenges in both capacitated and uncapacitated supply scenarios. The developed models demonstrate the feasibility and offer diverse interpretations of consumer choice behavior, drawing inspiration from established RUM-based frameworks. This research contributes to bridging the existing gap in the application of advanced DCMs within practical RM implementations.
@@ 1 Introduction
Advancements in Operations Research (OR) Modeling, coupled with innovative solving methods, have efficiently addressed large-scale challenges across economics, transportation, supply chain, revenue management, etc. The primary goal of these models is to optimize revenue, minimize costs, or mitigate pollution. Consequently, sophisticated demand models are under development to quantify these factors precisely. Recognizing the intricate interplay between demand and supply, our approach integrates both aspects, emphasizing simultaneous consideration of the supply side during supply chain planning. Notably, the pricing decision involves suppliers setting prices while buyers determine their willingness to pay, exemplified by the varying affordability of a \$10 rose in different neighborhoods. This underscores the crucial need for sellers to comprehend and anticipate customer purchasing behavior. Concurrently, Revenue Management (RM), within the realm of OR, concentrates on maximizing revenue through strategic pricing and capacity allocation. Our work uniquely combines economic and OR perspectives, incorporating customer behavior modeling for optimized supply systems.
Historically, demand and revenue-maximizing prices were determined using simple linear demand functions. Additionally, the main assumption of consumer demand was that a customer would buy a product if it were available and would not buy at all if it were not available [1]. Moreover, the supply and demand aspects were treated separately, meaning the supplier would decide on the price and capacity of the products regardless of the demand, which could be inefficient as the demand and supply sides are intricately intertwined [2].
@@
An essential assumption in RM is the variability of product prices for each customer. Precise understanding of customer behavior is necessary to optimize revenue. Discrete Choice Models (DCMs), which elucidate and predict customer choices, significantly contribute to customer behavior modeling. These models consider customer preferences, leading to efficient resource utilization and substantial revenue improvements. Numerous empirical studies underscore the impact of discrete choice modeling in diverse sectors such as RM [3], labor market [4], food industry [5], and airline industry [6].
@@
Random Utility Maximization Theory (RUM), established by [7, 8], forms the foundation for discrete choice modeling. The core assumption of RUM is that decision-makers are perfectly rational and aim to maximize their utilities. Recent studies, such as those by [9, 10, 11] propose an alternative approach by investigating DCMs derived from Random Regret Minimization (RRM) where the customer tries to minimize the regret caused by not choosing an alternative. Conditions exist where DCMs derived from either theory outperform the other, revealing diverse customer preferences [12, 13, 14].
@@
While existing literature often examines demand and supply separately or investigates supply based on a simplistic demand formulation, a comprehensive approach is needed for effective revenue generation [2]. Neglecting the interplay between supply and demand or focusing solely on one side can lead to suboptimal pricing strategies, resulting in reduced gains, the bullwhip effect, and discordance in supply chains. Recognizing that customer decisions drive demand underscores the vital importance of optimizing and coordinating both supply and demand concurrently. In this regard, [15] innovatively crafted a framework that characterizes how demand responds to suppliers’ decisions, linking supply and demand through price as a crucial variable. They linearized DCMs into predictors of demand using a Mixed Integer Linear Programming (MILP) model, emphasizing price variables in the objective function.
@@
In related literature, other works have incorporated DCMs in OR to forecast demand and maximize revenue. For example, [2, 16] addressed challenges arising from the non-linearity of DCM models in optimization. To address this, utility functions (RUM-based) were integrated into MILP constraints, introducing individual utility functions as scores to be maximized in discrete optimization, which is crucial for characterizing the demand side of the OR problem [17, 18].
To date, all demand estimation models in the realm of Revenue Management (RM) have relied on RUM-based DCMs to maximize revenue. To the best of our knowledge, there are no MILP models based on RRM theory to maximize revenue. In this paper, we introduce a novel MILP model developed based on RRM theory. The main distinction between our proposed algorithm and others lies in the underlying assumption about customer behavior. In MILP models based on RUM theory, it is assumed that customers are utility maximizers. These models determine prices based on the expectation that customers will maximize their utilities while suppliers focus on revenue maximization. Contrastingly, our algorithm operates on the premise that customers are regret minimizers. They select products aiming to minimize their regrets, while suppliers continue to strive for revenue maximization. This fundamental difference introduces novel challenges. In RUM-based MILP models, customers compute the expected utility gained from purchasing each product and select the one that maximizes their utility. Conversely, in our model based on RRM, we compare all products pairwise. This approach enables customers to choose the product with the least regret. To do so, we have designed the MILP model with additional constraints.
Existing literature in RM predominantly applies RUM-based MILP to maximize revenue and address the pricing issue. However, in this paper, an RRM-based MILP model has been proposed for the first time. The main contributions of this work are as follows:
  • 1.
    In many real-world scenarios, customers prioritize minimizing regret when making purchasing decisions, rendering RUM-based MILP models inadequate for accurate pricing and revenue maximization. To address this limitation, we have developed a MILP model based on an advanced RRM-based DCM. This model is capable of incorporating customers’ individual heterogeneity in revenue optimization and pricing strategies.
  • 2.
    Our proposed RRM-based model accommodates scenarios where suppliers have limited capacity to offer products to clients. Additionally, we have developed an alternative RRM-based model designed for scenarios where capacity is unlimited. There are two main challenges in developing such models:
    • (a)
      The DCMs become highly non-linear and non-convex when logit functions are used to estimate the probability of a customer choosing a product, as demonstrated by [2]. To overcome this challenge, we developed a linear formulation based on RRM theory, directly capturing how a customer selects a product to minimize regret. Additionally, we incorporated Gumbel-distributed noise to accommodate the uncertainties in consumer choices, which is consistent with RRM theory. Moreover, within the framework of our MILP model, we determine the average maximum revenue by sampling from the Gumbel distribution multiple times, a common practice in stochastic optimization approaches. This approach mirrors the technique used for the RUM-based MILP developed by [2].
    • (b)
      The computation of regret, even without considering the non-linearity introduced by DCMs, as mentioned in the previous point, inherently involves both minimization and maximization when comparing products. To handle this challenge, we employed linearization techniques in our MILP models for both capacitated and uncapacitated scenarios.
  • 3.
    We conducted experiments to assess the computational performance and solution quality of our proposed models compared to their RUM-based counterparts developed by [2, 15].
Finally, we demonstrate that when customers are regret minimizers, the maximum revenue achieved by our models differs significantly from that of RUM-based counterparts. This discrepancy can have a profound impact on the production plans of suppliers targeting regret-minimizing customers with their products. In particular, our results demonstrate that RUM-based MILP models can significantly cause a supplier to lose revenue when customers are regret minimizers.
The primary objective of this paper is to develop a robust framework for RM capable of maximizing revenue and optimizing product pricing for a supplier in scenarios with both capacitated and uncapacitated supply while accounting for customer choice behavior. Our hypothesis posits that customers are regret minimizers in certain scenarios, aiming to select products or opt out to minimize their regrets through pairwise comparisons of products. We contend that current frameworks relying on RUM theory may underestimate prices and lead to revenue losses. To address this, we consider both capacitated and uncapacitated supply quantities within our framework. Then, we introduce the RRM-based behavior of customers’ decision-making together with capacity, pricing, and revenue concerns to the MILP. Through experiments, we demonstrate that our framework based on RRM theory yields higher revenue by aligning with the behavior of customers under RRM theory.
For the dataset, we selected a range of numerical values representing potential product prices. Additionally, we incorporated randomly drawn samples from a Gumbel(0,1) distribution to represent the unobserved regrets and utilities of customers. To ensure robustness in our results, We generated multiple draws of such data, taking into account the stochastic nature of unobserved regrets and utilities, a methodology also employed by [2, 15].
Our results initially demonstrate the effectiveness of our developed models for both capacitated and uncapacitated versions. It is evident that the constrained version involves a larger number of variables and constraints compared to its unconstrained counterpart, resulting in longer solving times. Furthermore, our model solutions indicate that the RUM-based approach leads to revenue loss when customers are regret minimizers. This underscores the significance of employing RRM-based MILPs for such scenarios, which results in higher revenue for suppliers while still meeting customer needs. However, our models inherently entail more variables and constraints compared to their RUM-based counterparts, given the intrinsic nature of RRM theory. This characteristic inevitably leads to longer solution times.
@@ 2 Related Work
Within RM, pricing decisions hold prominent importance among various aspects to increase revenue. Several RM techniques aim to leverage customer heterogeneity for a product set at a specific time. They often focus on fixing alternatives or products and time dimensions while exploring customer differences. In our context, we emphasize the significance of customer heterogeneity in purchasing behavior. If all customers perceive a product uniformly with identical purchasing behavior, the customer dimension diminishes, leading to unexploited variations in willingness to pay. These variations, influenced by socio-economic factors like marital status, salary, age, gender, etc., play a critical role in maximizing revenue. On the other hand, optimizing revenue hinges on accurate demand prediction, which is crucial for effective supply planning and pricing strategies. The demand function is influenced by various variables, encompassing customer heterogeneity, prices, etc [19, 20]. Specifically, researchers have extensively focused on DCMs, particularly due to their capacity to predict demand by accounting for consumer heterogeneity in tastes and attributes [19, 8, 9]. RUM-based DCMs assume decision-makers are rational utility maximizers and models derived from these theories aim to reveal influential factors shaping customer choice preferences (e.g., see [21] for a recent review). However, in some cases, customers prefer regret minimization over utility maximization when making choices. Therefore, deriving DCMs from suitable behavioral theories is crucial [12, 22]. In this regard, [23] introduced a regret-based discrete choice model for consumer preference analysis with multinomial choice sets. Their empirical assessment, based on 43 comparisons between RUM and RRM models in the literature, revealed the superiority of RRM-based DCMs over RUM-based ones. The study demonstrated that both RRM-based and hybrid RRM-RUM-based DCMs consistently outperformed their RUM-based counterparts in terms of prediction accuracy and overall model fit. The authors emphasized that although performance differences were minor, the models led to significant managerial implications and diverse interpretations of behavior.
@@
The reference [24] conducted a choice experiment to investigate tourists’ hotel preferences. They created hypothetical hotel alternatives and randomized factors, including price, hotel rating, and walking time to key attractions, to form choice sets. RRM and RUM-based DCMs were utilized to estimate the coefficients of these factors, and they figured out the RRM-based DCMs performed significantly better than their RUM-based counterparts. Additionally, they showed that quantitative simulations can be performed to study the demand and pricing issues while incorporating customers’ heterogeneity into the model.
The reference [25] investigates the factors influencing park-and-ride (PNR) lot choice behavior among users. Two models, based on RUM and RRM concepts, are developed to understand why users select one PNR lot over another. Travel time in both the auto and transit networks is considered in the models, with data sourced from strategic transport networks and Google’s transit feed specification. The study explores the trade-off relationship between the two networks and suggests that RRM models may better capture this compromise effect compared to traditional RUM models. Results show that RRM models, particularly the classical RRM, provide insights into transport choice behavior and offer improved predictive abilities compared to RUM models. Additionally, this study contributes to the understanding of RRM models by estimating them on revealed preference data.
The reference [26] examines the use of RRM models in analyzing public preferences for air quality improvement policies, compared to traditional RUM models. The research finds that RRM models outperform RUM models in terms of goodness of fit and predictive accuracy. Approximately 55.9 % of respondents’ choices are influenced by regret. The study also reveals differences in preferences between regret-driven and utility-driven respondents: regret-driven individuals prioritize increased clean air days, reduced haze days, and lower mortality, while utility-driven individuals prioritize shortening policy delay. These findings can inform the design of socially optimal air quality policies and contribute to the development of environmental policies.
The reference [27] addresses the complex factors influencing route choice behavior in the context of advancing traffic management systems and the availability of diverse traffic information sources. While previous studies have focused on travel time and cost, this study examines the impact of information provision mechanisms on route preferences. Conducted in the Greater Orlando Region, USA, the research utilizes web-based stated preference surveys to gather data. The analysis employs RUM and RRM models to quantify the trade-offs among different attributes affecting route choice. Various variables, including trip characteristics, roadway features, and traffic information attributes, are incorporated into the models. The results offer insights into the trade-offs involved in travel information provision, aiding in the design of effective active traffic management systems. Additionally, the research customizes the computation for RRM models to account for variable interactions, enhancing the understanding of route choice behavior.
The reference [28] investigates whether RRM models provide a better description of evacuation behavior compared to traditional RUM models. The researchers designed a revealed preference (RP) survey focusing on evacuation choices in the aftermath of the 2017 Southern California Wildfires. Contrary to their hypothesis, the data did not support the superiority of RRM over RUM for evacuation decisions. The lack of sufficient variation in attribute levels across alternatives made it challenging to distinguish between linear utility and non-linear regret. However, the study identified weak regret aversion for certain attributes and detected class-specific regret for route and mode choices using a mixed-decision rule latent class choice model. Despite the initial findings, the researchers suggest that RRM for evacuations could still be promising with further exploration. The study highlights methodological implications for RP studies with complex choice sets and limited attribute variability.
@@
Additionally, DCM models can be considered as machine learning models. For instance, multinomial logit models correspond to well-known logistic regression models in ML. Machine algorithms can be broadly categorized into one of the following three types: supervised, unsupervised, and agent-based learning, with DCMs falling into the supervised category. Paper [29] provides a comprehensive overview of the existing ML techniques along with their types, characteristics, and potential applications. Indeed, ML tools such as MNL and MMNL are used to predict customers’ choices based on RUM theory [30, 31, 32, 33, 34]. Reinforcement learning as an online learning method can also be leveraged to help RM [35, 36, 37, 38, 39].
Accurately predicting demand is a crucial aspect of revenue optimization, vital for effective supply planning and pricing strategies. Demand forecasting relies on various factors, such as customer diversity, pricing, competitor offerings, brand reputation, and product attributes. Traditionally, demand estimation employed a simple linear function, focusing solely on product price, which proved to be inadequate as it overlooked many influential factors. Additionally, the conventional assumption that customers faced only binary choices — either the product was available for purchase or not due to capacity constraints — did not reflect the complexities of real-world consumer behavior [1]. Recognizing these limitations, segmentation of customer offers emerged in the early 21st century, categorizing customers based on demographics like age, income, etc., and providing tailored choice sets and prices accordingly [3, 1]. Confronted with limited capacity, stochastic dynamic programming models emerged to define product offer sets while accounting for capacity constraints over selling periods. Despite their efficacy, these models encountered difficulties in addressing large-scale problems [1]. Moreover, they overlooked customer individual-level heterogeneity, typically categorizing customers into broad groups. The reference [40] made a pioneering contribution to Revenue Management by focusing on modeling passenger preferences using Discrete Choice Models (DCMs). They considered the heterogeneity of customers and estimated the probability of purchasing a product using DCMs, integrating it into their stochastic dynamic programming model. Following [40], [41] addressed a general question in RM related to consumer choice behavior (RUM-based). They explored the network revenue management problem with flexible products, considering scenarios where demand was either exogenously given or generated by a consumer choice model based on RUM-based DCM. Their findings indicated that the optimal value of the stochastic optimization problem closely approximated that of a deterministic linear program. The reference [42] delved into the implications of the concept introduced by [41] in RM. They developed a Choice-based Deterministic Linear Programming (CDLP) model under consumer choice behavior, focusing on providing suitable offer sets, seats, and flights in the airline industry while adhering to capacity constraints over time. Their study demonstrated that the CDLP, inspired by [41], serves as a deterministic approximation of the stochastic formulation, with its solution being an asymptotic optimal solution for the original stochastic problem influenced by the stochasticity of customer choice probability. The reference [43] expanded on the CDLP introduced by [42]. They addressed the issue of overlapping offer sets in the airline industry, where airlines typically present affordable morning tickets in one set and pricier afternoon tickets in another set. This extension considered combinations such as expensive and morning tickets, providing insights into more realistic scenarios. They applied the column generation method to tackle the problem’s dimensionality issue. Particularly, meta-heauristc methods seem to be a promising approach in approximating high-quality solutions for large-sized problems (e.g., see [44, 45, 46, 47, 48]). The reference [49] introduced a modified CDLP model, solving it on a synthetic dataset to obtain optimal offer sets and an upper bound on revenue. Additionally, they suggested a non-parametric model for estimating choice probabilities, utilizing the synthetic data employed in the CDLP.
Thus far, we have elucidated the role of certain OR techniques, such as Stochastic programming and Linear programming, in RM. However, despite some models considering individual customer heterogeneity and RUM-based DCMs, they have typically assumed a fixed price for all customers or customer segments. This overlooks the influence of price dynamics on customer choice and demand, as demand is directly influenced by customer decisions. To address this limitation, the pricing of products should be treated as a variable rather than a fixed parameter in the models. This challenge was first tackled by [50], who introduced product prices as variables alongside considerations of individual customer heterogeneity and capacity constraints. Consequently, conventional RUM-based DCMs, which previously predicted the probability of a customer purchasing a product at a fixed price, transformed into nonlinear and non-convex demand functions due to the inclusion of price as a variable. To mitigate this complexity, [50] linearized the RUM-based DCMs using MILP models, thereby pioneering the application of RUM-based MILP models in RM. Subsequently, [15] extended this approach by linearizing a mixed multinomial logit DCM, which considers unobserved customer utility correlated with customer attributes, through MILP techniques. Their objective was to maximize revenue by simultaneously optimizing supply (price and capacity) and demand (price, utility levels, customer satisfaction) factors. The reference [51] delved into an economics-oriented approach, emphasizing passenger satisfaction maximization by integrating DCMs into MILP models. The objective was to enhance the public transportation network by utilizing toll revenue from highways, ensuring a balance between economic considerations and passenger contentment. To date, no RRM-based MILP model has been applied for in RM.
@@
In conclusion, DCMs are state-of-the-art for estimating dis-aggregate demand, showing promise in RM. The literature demonstrates integrating DCMs into optimization models for better supply-demand interactions. However, a gap exists between DCMs derived from theories like RRM and optimization formulations. Finally, [2, 15, 52, 53] developed a RUM-based DCM framework where customers are utility maximizers. We extended the framework for RRM-based DCM introduced by [10], where the customers are regret minimizers.